322 L of hydrogen occupies a volume of 197 L at STP. If the initial temperature of the hydrogen was 37 C, what was its initial pressure?


1
Expert's answer
2022-02-27T17:01:05-0500

Using ideal gas law

p = nRT/V

Density (H2) = 0.000089 g/ml

Mass (H2) = Volume (H2) x Density (H2) = 322000 ml x 0.000089 g/ml = 28.658 g

n(H2) = m(H2) / Mr(H2) = 28.658 g / 2 g/mol = 14.32 moles

p = 14.32 moles x 0.08206 L atm / mol K x (37 + 273 K) / 197 L = 1.84 atm